Following are several figures reported for Allister and Barone as of December 31, 2018:

Allister Barone
Inventory $ 410,000 $ 210,000
Sales 820,000 620,000
Investment income not given
Cost of goods sold 410,000 310,000
Operating expenses 185,000 255,000

Allister acquired 80 percent of Barone in January 2017. In allocating the newly acquired subsidiary's fair value at the acquisition date, Allister noted that Barone had developed a customer list worth $60,000 that was unrecorded on its accounting records and had a 4-year remaining life. Any remaining excess fair value over Barone's book value was attributed to goodwill. During 2018, Barone sells inventory costing $121,000 to Allister for $162,000. Of this amount, 10 percent remains unsold in Allister's warehouse at year-end.

Determine balances for the following items that would appear on Allister's consolidated financial statements for 2018:

INVENTORY

SALES

COST OF GOOD SOLD

OPERATING EXPENSES

NET INCOME ATTRIBUTABLE TO NONCONTROLLING INTEREST
Answer #1
Customer list amortization =$60000/4 years $15000 per year
Intra-entity gross profit =$162000-121000 $41000
Inventory remaining at year end 10%
Intra-entity gross profit in ending inventory,12/31 =$41000*10% $4100
Consolidated totals:
Inventory =410000+210000-4100 $615900
(add the two book values and subtract the ending intra-entity gross profit of $4100)
Sales =820000+620000-162000 $1278000
(add the two book values and subtract the $162000 intra-entity transfer)
Cost of goods sold =410000+310000-162000+4100 $562100
(add the two book values and subtract the intra-entity transfer and add [to defer] ending inventory intra-entity gross profit)
Operating expenses =185000+255000+15000 $455000
(add the two book values and the amortization expense for the period)
Net income attribute to non controlling intersest =10%*(620000-310000-255000-4100-15000) $3590

(Gross profit deferral is allocated to the noncontrolling interest because the transfer was upstream from Barone to Allister.)